Fermi National Laboratory
 

Accelerator Update for February 16, 2007 - February 19, 2007

 

Friday February 16

The day shift began with Operations monitoring the Recycler stash (213.2E10) and the Antiproton Source (Pbar) stack (48.6E10), with the Booster, Linac, Main Injector (MI), and Recycler operating normally, with MiniBooNE off, with NuMI taking beam, with cryo system technicians working on the wet engine at Tevatron (TeV) sector A1, with D-Zero (D0) in access, and with the Collider Detector at Fermilab (CDF) waiting for the colliding beams.

Operations reported at 8:45 AM that power supply experts completed their work on the TeV sector A1 bulk power supplies; they had been drenched by an LCW leak.

D0 came out of access at 9:22 AM.

Cryo system techs reported at 10:39 AM that they had completed their work on the TeV sector A1 wet engine.

At 11:37 AM, the MI-20 BPMs stopped responding. An MI expert investigated.

Operations ramped the TeV at 12:34 PM.

At 12:35 PM, a Booster RF station (BRF12) tripped off. It reset.

Operations put the TeV through a dry squeeze at 12:49 PM.

Operations put the TeV into shot setup at 1:36 PM.

Recycler experts began a RF study at 3:46 PM.

Operations established store 5231 at 3:59 PM with an initial luminosity of 246.83E30.

Operations reported around 4:20 PM that a Linac RF station (LRF3) had tripped off. Operators discovered a problem with the controls for its LCW water skid. Operations contacted experts.

Linac experts replaced a 24-volt power supply on the LRF3 water skid. NuMI and Pbar resumed taking beam at 6:41 PM.

At 9:17 PM, the 750keV area gradient buncher dropped to zero and wouldn’t remotely come up. Operators investigated and had it back up within 15 minutes.

Operations reported at midnight that the Recycler had stashed antiprotons three times during the evening shift.

Saturday February 17

The midnight shift began with Operations monitoring store 5231 (87.64E30) and the stash (103.2E10), with Pbar stacking (57.0E10), and with NuMI taking beam.

At 12:30 AM, BRF15 & 16 tripped off when a vacuum pump tripped off. Everything reset.

Around 5 AM, the Pbar Electron Cooling tripped off due to its MPS modulation going to zero. Operations contacted experts and then set the modulation to its normal amplitude.

At 7:30 AM, a Linac Klystron RF station (KRRF3) tripped off on a LCW flow indication. Operators investigated and reset the station.

At 7:58 AM, Operations tried to turn on the Meson power supplies, but a problem with a special M01 LCW skid was holding off the permits. Operations contacted experts.

Operations reported at 8 AM that the Recycler had stashed three times during the midnight shift.

Operations established beam to Switchyard (SY) and Meson at 10:52 AM.

At 10:57 AM, KRF5 tripped off for the first of two times in two minutes. It reset both times.

At 11:56 AM, Operations reported that a control panel issue with a power supply (H201) had halted beam to SY. Operations contacted experts.

At 12:41 PM, store 5231 quenched at TeV sectors B0 and B1. Status indicated a sector B2 power supply failure. Further investigation showed that a quadrupole power supply (C:B0Q6) had tripped just before the quench. Operations called in experts.

Power supply experts reported that they had replaced a bridge diode for H201. SY beam resumed at 2:20 PM.

At 4:14 PM, H201 tripped off again. Operators investigated and found that the control power breaker had again tripped off. Operations contacted an expert.

Operations reported at 7:55 PM that the Recycler had stashed four times since 8 AM.

Operations reported at 8 PM that the Pbar stack was at 32.9E10, the stash was at 332.7E10, and that the TeV sector B0 cryo system was recovering.

At 9:33 PM, Operations reported that a Recycler transfer had bombed out due to an incrementing problem. Operators completed the transfer manually. This happened several more times. Operations talked to experts and between them discovered a bad setting.

Operations put the TeV through a dry squeeze at 10:21 PM.

At 10:37 PM, a TeV quadrupole power supply (D0Q5) tripped. Operations contacted an expert. At 11:46 PM, with help from experts, operations determined that the fault was due to a magnet lead failure.

A power supply expert completed work on H201 at 11 PM.

Sunday February 18

The midnight shift began with the TeV recovering from a quench, with Operations monitoring the stash, with Pbar stacking, and with NuMI taking beam.

Operations put the TeV back into a dry squeeze at 12:01 AM.

At 12:39 AM, Operations reported a problem with the TeV sector D4 BPMs not reading back. Thier VME was not responding. Rebooting the VME didn’t help. Operations tried to contact a VME expert for help, but couldn’t find one. A TeV expert reported that Operations could run the TeV without these BPMs.

Operations put the TeV into a dry squeeze at 1:07 AM.

Operations put the TeV into shot setup at 1:37 AM.

Operations established store 5234 at 3:21 AM with an initial luminosity of 285.03E30. This store set a New Record Luminosity.

At 3:49 AM, Operations reported some beam fall out during a Recycler stash. Operators talked to experts and plotted devices, but could not discover the problem.

Operations reported at 8:25 AM that the Recycler had stashed three times during the midnight shift.

At 8:50 AM, at Pbar Debuncher power supply (D:H704) tripped off on a phase imbalance indication. It would not reset. Operations turned off beam to Pbar and investigated. Operations called in an expert.

At 9:31 AM, KRF4 tripped off. It reset.

Operations established beam to SY at 9:44 AM.

At 11:08 AM, LRF3 tripped off twice in 30 minutes. It reset both times.

Power supply experts discovered and replaced a bad 15-volt supply on D:H704.

At 4:04 PM, Operations reported that the Pbar antiproton production had dropped. Operators found that a tuning program had disrupted the beam in AP1. Operations restored settings and restarted a tuning program — normal antiproton production resumed.

SY/Meson/MTest experts reported at 4:28 PM that they were done with the beam for the day.

Operations reported at 8 PM that the Recycler had stashed three times during the evening shift.

Operations reported the following status at 8 PM:

  • Store 5234 luminosity = 62.08E30
  • Stash = 218.2E10
  • Stack = 27.0E10

At 10 PM, Operations reported control issues with the MI-60 RF vacuum permits. Half of the permits showed bad status even though the pumps were on. Operations decided to contact an expert in the morning.

Monday February 19

The midnight shift began with Operations monitoring store 5234 and the stash, with Pbar stacking, and with NuMI taking beam.

At 12:28 AM, BRF6 tripped off. It reset okay.

At 4:23 AM, BRF6 tripped off again. It reset.

TeV experts began at 6:03 AM to use the Tevatron Electron Lens (TEL2) to clean the abort gap, and then began an end-of-store study and a crystal collimator study.

Operations reported at 7:57 AM that the Recycler had stashed five times since 8 PM.

Operations terminated store 5234 at 8:04 AM.

The Plans for Monday

The plans for today are to put the TeV into shot setup this morning. After the shot, the Run Coordinator will allow an access into Booster and MI for maintenance work.


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 02/19/2007    email Fermilab